OpenStudy (anonymous):

F(x) varies inversely with x and f(x) = –15 when x = 30. What is the inverse variation equation? a. f(x) = -5/x b. f(x) = -2/x c. f(x) = -90/x d. f(x) = -450/x

OpenStudy (anonymous):

Wait i think it is D.

OpenStudy (anonymous):

f(x) = -15, x=30 -15 = k/30 k = (-15)(30) k = -450 f(x) = -450/x right?

OpenStudy (anonymous):

absolutely correct :)